A semi-circle of radius 14 cm is formed from a piece of wire. If it is bent into a rectangle whose length is 2 cm more than its width, find the area of the rectangle.


1
Expert's answer
2021-05-17T03:55:02-0400

"\text{
perimeter of semi circle}=(2+\pi)r=(2+\pi)×14=71.9823 cm\newline\text{
Given, length of rectangle(l) is 2cm more than its width(b).}\newline
l=2+b\newline\text{
perimeter of rectangle}=2(l+b)\newline\text{
Given, perimeter of semi circle=perimeter of rectangle}\newline
(2+\pi)×14=2(l+2+l)\newline
(2+\pi)×14=4(l+1)\newline
l+1=\frac{(2+\pi)×14}{4}\newline
l=\frac{(2+\pi)×14}{4}-1\newline
l=16.9956 cm\newline\text{
Therefore, width of rectangle is }18.9956 cm.\newline
\text{Area of the rectangle}=lb=16.9956×18.9956=322.8416 cm^2"